# Towards Backdoor Stealthiness in Model Parameter Space

### v2

@@ 6. Conclusions & Future Work
Full version with appendix: (Xu et al., 2025).
@@ A.1. Datasets
CIFAR10. The CIFAR10 (Krizhevsky et al., 2009) contains 50,000 training images and 10,000 testing images with the size of $3\times 32\times 32$ in 10 classes.
GTSRB. The GTSRB (Stallkamp et al., 2012) contains 39,209 training images and 12,630 testing images in 43 classes. In our experiments, the images are resized to $3\times 32\times 32$.
ImageNet200. ImageNet (Deng et al., 2009) contains over 1.2 million high-resolution images in 1,000 classes. In our experiments, we randomly select 200 classes from the ImageNet dataset as our ImageNet200 dataset. Each class has 1,300 training images and 50 testing images. The ImageNet images are resized to $3\times 224\times 224$.
@@ A.2. Backdoor Attacks
Our attack is compared with 12 well-known and representative attacks: BadNets (Gu et al., 2019), Blend (Chen et al., 2017), WaNet (Nguyen and Tran, 2021), IAD (Nguyen and Tran, 2020), AdvDoor (Zhang et al., 2021), BppAttack (Wang et al., 2022b), LC (Turner et al., 2019), Narcissus (Zeng et al., 2023), Adap-Blend (Qi et al., 2023a), SSDT (Mo et al., 2024), DFST (Cheng et al., 2021), and DFBA (Cao et al., 2024).
@@ A.3. Backdoor Defenses
We evaluate our attack and baseline attacks against 17 defenses, including 4 pruning-based methods (FP (Liu et al., 2018a), ANP (Wu and Wang, 2021), CLP (Zheng et al., 2022), and RNP (Li et al., 2023a)), 5 fine-tuning-based methods (vanilla FT, FT-SAM (Zhu et al., 2023), I-BAU (Zeng et al., 2022), FST (Min et al., 2023), and BTI-DBF(U) (Xu et al., 2024a)), 5 backdoor model detections (NC (Wang et al., 2019), Tabor (Guo et al., 2020), FeatureRE (Wang et al., 2022a), Unicorn (Wang et al., 2023), and BTI-DBF (Xu et al., 2024a)), 2 backdoor input detections (Scale-up (Guo et al., 2023) and IBD-PSC (Hou et al., 2024)), and a proactive detection CT (Qi et al., 2023b).
FP [^9] https://github.com/SCLBD/BackdoorBench/blob/main/defense/fp.py, vanilla FT [^10] https://github.com/SCLBD/BackdoorBench/blob/main/defense/ft.py, FT-SAM [^11] https://github.com/SCLBD/BackdoorBench/blob/main/defense/ft-sam.py, I-BAU [^12] https://github.com/SCLBD/BackdoorBench/blob/main/defense/i-bau.py. We use the implementation and default hyperparameters from BackdoorBench (Wu et al., 2022). For FT-SAM on ImageNet200, the default setting will decrease benign accuracy to 0.465, so we reduce its training schedule to 25 epochs. Please note that the experiments on CIFAR10 with FT-SAM usually converge within 20 epochs in our experiments. Thus, decreasing the training schedule is not harmful to the defense performance.
NC [^13] https://github.com/ain-soph/trojanzoo/blob/main/trojanvision/defenses/backdoor/model_inspection/neural_cleanse.py and Tabor [^14] https://github.com/ain-soph/trojanzoo/blob/main/trojanvision/defenses/backdoor/model_inspection/tabor.py. We use the implementation from TrojanZoo (Pang et al., 2022). 1% training set and 100 epochs are used for trigger inversion.
Scale-up [^15] https://github.com/THUYimingLi/BackdoorBox/blob/main/core/defenses/SCALE_UP.py, IBD-PSC [^16] https://github.com/THUYimingLi/BackdoorBox/blob/main/core/defenses/IBD_PSC.py. We use the implementation and default hyperparameters from BackdoorBox (Li et al., 2023b).
@@ Appendix B Attack Summary
BadNets (Gu et al., 2019)
Blend (Chen et al., 2017)
WaNet (Nguyen and Tran, 2021)
IAD (Nguyen and Tran, 2020)
AdvDoor (Zhang et al., 2021)
Bpp (Wang et al., 2022b)
LC (Turner et al., 2019)
Narcissus (Zeng et al., 2023)
SSDT (Mo et al., 2024)
Adap-blend (Qi et al., 2023a)
DFST (Cheng et al., 2021)
DFBA (Cao et al., 2024)
@@
Attack Scale-up (Guo et al., 2023) IBD-PSC (Hou et al., 2024)
@@
BadNets (Gu et al., 2019) 81.93 32.90 0.7627 0.7524 100 7.90 0.9996 0.9606
Blend (Chen et al., 2017) 99.32 38.74 0.8681 0.8275 100 0.90 1.00 0.9953
Adap-Blend (Qi et al., 2023a) 68.72 18.99 0.7621 0.7297 53.95 11.77 0.8731 0.6495
@@
Victim Surrogate No Defense FT-SAM (Zhu et al., 2023) I-BAU (Zeng et al., 2022) FST (Min et al., 2023)
@@ Appendix C Additional Experiments
C.1. Detection of backdoor input.
Backdoor input detection is a defense technique that determines whether or not a given input includes a backdoor trigger. We show that Grond-generated backdoor samples can resist established backdoor detection methods. Table 14 shows the input-space detection results using Scale-up (Guo et al., 2023) and IBD-PSC (Hou et al., 2024). We report the True Positive Rate (TPR), False Positive Rate (FPR), AUC, and F1 score in Table 14 for baseline attacks and Grond, where Scale-up and IBD-PSC are effective against three baseline attacks but cannot detect Grond-generated backdoor samples.
@@ C.2. Different Architectures with Different Surrogate Models
We evaluate Grond with four additional victim architectures in Table 15: VGG16, DenseNet121, EfficientNet-B0, and InceptionNeXt-Tiny. In addition, as Grond requires a surrogate model to generate UPGD as the backdoor trigger, we provide the results when UPGD is generated using different architectures for the surrogate model. For each architecture, UPGD is generated by either the victim architecture or ResNet18 to perform our attack. In Table 15, we use the three most powerful defenses according to Tables 2 and 3. Regardless of the model’s architectures or the architectures for UPGD, Grond bypasses most defenses. This is because the UPGD contains semantic information of the target class and can be transferred among different architectures (Moosavi-Dezfooli et al., 2017). In a few cases, using UPGD generated by the same architecture shows better attack performance. For example, conducting Grond on InceptionNeXt-Tiny with UPGD generated by InceptionNeXt-Tiny shows ASRs above 90%, but also a much lower ASR when using UPGD generated by ResNet18. We conjecture that transferring UPGD from ResNet18 to InceptionNeXt-Tiny is more difficult than transferring it to other architectures due to the large convolution kernel design of InceptionNeXt.
@@ C.3. Large and Transformer-Based Architectures
To further evaluate the effectiveness of Grond on large-sized and especially transformer-based models, we conduct experiments on various architectures, including ResNet101 (He et al., 2016), Swin-Base (Liu et al., 2021), Swin-Large (Liu et al., 2021), ViT-Base (Dosovitskiy et al., 2021), and ViT-Large (Dosovitskiy et al., 2021).
There are two components (UPGD trigger and ABI) of Grond that are necessary to be applied to transformer-based architectures. The UPGD trigger generation algorithm is compatible with transformer-based architectures, as transformers’ gradient information can be used to generate UPGD perturbations. However, UCLC (Zheng et al., 2022) for ABI is incompatible with the attention mechanism. The UCLC is designed for feature channels of convolutional layers, which do not exist in the attention mechanism. Therefore, we design a tailored novel pruning method against backdoor neurons in the attention-based models. Specifically, we consider an attention head as a “neuron”, the new pruning method then calculates the singular value decomposition of each attention head as an index of relevance to the backdoor effect. The process can also be considered as solving for the approximation of the Lipschitz Constant for each attention head according to (Zheng et al., 2022). Then, the sensitive heads are updated using the following equation:
(4) ${\bm{\theta}}_{l}^{(k)}:=\left\{\begin{aligned} &{\bm{\theta}}_{l}^{(k)}\times 0.5,&\sigma({\bm{\theta}}_{l}^{(k)})>\text{mean}(\sigma({\bm{\theta}}_{l}))+u\times\text{std}(\sigma({\bm{\theta}}_{l}))\\ &{\bm{\theta}}_{l}^{(k)},&\text{otherwise,}\end{aligned}\right.$
@@
Dataset Victim Surrogate Parameters (M) No Defense vanilla FT FT-SAM (Zhu et al., 2023) FST (Min et al., 2023)
@@ C.4. Backdoor analysis by feature space inversion
We also provide a feature space analysis for different attacks by using a feature mask to decouple the benign and backdoor features following BTI-DBF (Xu et al., 2024a) and BAN (Xu et al., 2024b). The decoupling assumes that benign features related to the correct prediction introduce a lower loss, and the backdoor features related to backdoor prediction introduce a higher loss. In particular, the benign and inversed backdoor features are decoupled as follows:
@@
where $g=f_{L-1}\circ\cdots f_{1}$ is the model without the classification head. ${\bm{m}}$ is the learned feature mask, and $\mathcal{D}_{l}$ is a small set of benign samples with correct labels. As validated with BTI-DBF (Xu et al., 2024a) and BAN (Xu et al., 2024b), benign and backdoor features can be decoupled by the mask ${\bm{m}}$, after which backdoor features will introduce a substantially higher loss with respect to the ground truth label.
@@
Hyperparameters for the Inversed Backdoor Feature Loss Following the settings in BTI-DBF (Xu et al., 2024a) and BAN (Xu et al., 2024b), we use Adam and the learning rate of 0.01 to search for 20 epochs for the feature mask in Equation 5. The optimization of the mask uses 1% of the training data. $\lambda$ is 0.72. The elements in the mask are limited to continuous values between 0 and 1.
@@
Dirty-Label Grond PR No Defense CLP (Zheng et al., 2022) FT-SAM (Zhu et al., 2023)
@@
C.5. Dirty-label Grond.
@@
C.6. Adversarial Backdoor Injection Does Not Impact Backdoor Effectiveness in Case of No Defense.
@@ C.7. Further TAC analysis
Table 18. Evaluation of Grond against adversarial training (AT) on CIFAR10. RA refers to accuracy under PGD (Madry et al., 2018) adversarial attack.
Arch No Defense AT
@@
Figure 12. The weight changes after FT-SAM (Zhu et al., 2023) for baseline attacks.
@@ C.9. Weight Changes After FT-SAM with 10 Baseline Attacks
Figure 12 provides the results of weight changes after applying FT-SAM (Zhu et al., 2023) to ten baseline attacks. The weight changes in baseline attacks show a significant difference in a few prominent neurons.
@@ C.10. Attempts to Craft Inputs that Minimize Changes in the Parameter Space
In the previous design of Grond, the backdoor is injected by the UPGD trigger. Then, Grond, through ABI, limits the changes of backdoor neurons by reducing their weights. We consider replacing the ABI component with an operation only in the training data, which does not require control over training. Our goal is to partially break the effectiveness of injecting the backdoor via the poisoned data. Specifically, we add poisoned data into the target class with a poisoning rate of $PR=5\%$ under the clean-label setting, following the standard setting in previous work (Wang et al., 2023; Xu et al., 2024b). Then, we add poisoned data into every class with a poisoning rate of $PR\times\lambda$. We aim to use the poisoned data in non-targeted classes to partially break the effectiveness of poisoned data in the targeted class. Table 19 provides the experimental results with different $\lambda$ values. With the decrease in $\lambda$, both ASR (no defense) and ASR (after applying CLP) increase. This means that our attempt to minimize changes in the parameter space for stealthiness works. However, ASR after applying CLP is still lower than 10%, which means the attack is not robust against the backdoor defense.

### v3

@@ 6. Conclusions & Future Work
Full version with appendix: (xu2025towards).
@@ A.1. Datasets
CIFAR10. The CIFAR10 (krizhevsky2009cifar10) contains 50,000 training images and 10,000 testing images with the size of $3\times 32\times 32$ in 10 classes.
GTSRB. The GTSRB (Stallkamp2012gtsrb) contains 39,209 training images and 12,630 testing images in 43 classes. In our experiments, the images are resized to $3\times 32\times 32$.
ImageNet200. ImageNet (jia2009imagenet) contains over 1.2 million high-resolution images in 1,000 classes. In our experiments, we randomly select 200 classes from the ImageNet dataset as our ImageNet200 dataset. Each class has 1,300 training images and 50 testing images. The ImageNet images are resized to $3\times 224\times 224$.
@@ A.2. Backdoor Attacks
Our attack is compared with 12 well-known and representative attacks: BadNets (Gu_2019_badnet), Blend (chen_2017_blend), WaNet (nguyen2021wanet), IAD (nguyen2020inputaware), AdvDoor (zhang2021advdoor), BppAttack (Wang_2022_bpp), LC (turner2019lc), Narcissus (zeng2023narcisus), Adap-Blend (qi2022revisitingadaptive), SSDT (mo2024ssdt), DFST (Cheng2021dfst), and DFBA (cao2024dfba).
@@ A.3. Backdoor Defenses
We evaluate our attack and baseline attacks against 17 defenses, including 4 pruning-based methods (FP (liu2018finepruning), ANP (Wu_2021_anp), CLP (zheng2022clp), and RNP (li2023rnp)), 5 fine-tuning-based methods (vanilla FT, FT-SAM (Zhu2023ftsam), I-BAU (zeng2022ibau), FST (min2023fst), and BTI-DBF(U) (xu2024btidbf)), 5 backdoor model detections (NC (wang_2019_nc), Tabor (guo_2020_tabor), FeatureRE (wang2022featurere), Unicorn (wang2023unicorn), and BTI-DBF (xu2024btidbf)), 2 backdoor input detections (Scale-up (guo2023scaleup) and IBD-PSC (hou2024ibdpsc)), and a proactive detection CT (qi2023proactivedetection).
FP [^9] https://github.com/SCLBD/BackdoorBench/blob/main/defense/fp.py, vanilla FT [^10] https://github.com/SCLBD/BackdoorBench/blob/main/defense/ft.py, FT-SAM [^11] https://github.com/SCLBD/BackdoorBench/blob/main/defense/ft-sam.py, I-BAU [^12] https://github.com/SCLBD/BackdoorBench/blob/main/defense/i-bau.py. We use the implementation and default hyperparameters from BackdoorBench (wu2022backdoorbench). For FT-SAM on ImageNet200, the default setting will decrease benign accuracy to 0.465, so we reduce its training schedule to 25 epochs. Please note that the experiments on CIFAR10 with FT-SAM usually converge within 20 epochs in our experiments. Thus, decreasing the training schedule is not harmful to the defense performance.
NC [^13] https://github.com/ain-soph/trojanzoo/blob/main/trojanvision/defenses/backdoor/model_inspection/neural_cleanse.py and Tabor [^14] https://github.com/ain-soph/trojanzoo/blob/main/trojanvision/defenses/backdoor/model_inspection/tabor.py. We use the implementation from TrojanZoo (pang_2022_trojanzoo). 1% training set and 100 epochs are used for trigger inversion.
Scale-up [^15] https://github.com/THUYimingLi/BackdoorBox/blob/main/core/defenses/SCALE_UP.py, IBD-PSC [^16] https://github.com/THUYimingLi/BackdoorBox/blob/main/core/defenses/IBD_PSC.py. We use the implementation and default hyperparameters from BackdoorBox (li2023backdoorbox).
@@ Appendix B Attack Summary
BadNets (Gu_2019_badnet)
Blend (chen_2017_blend)
WaNet (nguyen2021wanet)
IAD (nguyen2020inputaware)
AdvDoor (zhang2021advdoor)
Bpp (Wang_2022_bpp)
LC (turner2019lc)
Narcissus (zeng2023narcisus)
SSDT (mo2024ssdt)
Adap-blend (qi2022revisitingadaptive)
DFST (Cheng2021dfst)
DFBA (cao2024dfba)
@@
Attack Scale-up (guo2023scaleup) IBD-PSC (hou2024ibdpsc)
@@
BadNets (Gu_2019_badnet) 81.93 32.90 0.7627 0.7524 100 7.90 0.9996 0.9606
Blend (chen_2017_blend) 99.32 38.74 0.8681 0.8275 100 0.90 1.00 0.9953
Adap-Blend (qi2022revisitingadaptive) 68.72 18.99 0.7621 0.7297 53.95 11.77 0.8731 0.6495
@@
Victim Surrogate No Defense FT-SAM (Zhu2023ftsam) I-BAU (zeng2022ibau) FST (min2023fst)
@@ Appendix C Additional Experiments
C.1. Detection of backdoor input
Backdoor input detection is a defense technique that determines whether or not a given input includes a backdoor trigger. We show that Grond-generated backdoor samples can resist established backdoor detection methods. Table 14 shows the input-space detection results using Scale-up (guo2023scaleup) and IBD-PSC (hou2024ibdpsc). We report the True Positive Rate (TPR), False Positive Rate (FPR), AUC, and F1 score in Table 14 for baseline attacks and Grond, where Scale-up and IBD-PSC are effective against three baseline attacks but cannot detect Grond-generated backdoor samples.
@@ C.2. Different Architectures with Different Surrogate Models
We evaluate Grond with four additional victim architectures in Table 15: VGG16, DenseNet121, EfficientNet-B0, and InceptionNeXt-Tiny. In addition, as Grond requires a surrogate model to generate UPGD as the backdoor trigger, we provide the results when UPGD is generated using different architectures for the surrogate model. For each architecture, UPGD is generated by either the victim architecture or ResNet18 to perform our attack. In Table 15, we use the three most powerful defenses according to Tables 2 and 3. Regardless of the model’s architectures or the architectures for UPGD, Grond bypasses most defenses. This is because the UPGD contains semantic information of the target class and can be transferred among different architectures (Moosavi2017uap). In a few cases, using UPGD generated by the same architecture shows better attack performance. For example, conducting Grond on InceptionNeXt-Tiny with UPGD generated by InceptionNeXt-Tiny shows ASRs above 90%, but also a much lower ASR when using UPGD generated by ResNet18. We conjecture that transferring UPGD from ResNet18 to InceptionNeXt-Tiny is more difficult than transferring it to other architectures due to the large convolution kernel design of InceptionNeXt.
@@ C.3. Large and Transformer-Based Architectures
To further evaluate the effectiveness of Grond on large-sized and especially transformer-based models, we conduct experiments on various architectures, including ResNet101 (He_2016_resnet), Swin-Base (Liu2021swin), Swin-Large (Liu2021swin), ViT-Base (dosovitskiy2021vit), and ViT-Large (dosovitskiy2021vit).
There are two components (UPGD trigger and ABI) of Grond that are necessary to be applied to transformer-based architectures. The UPGD trigger generation algorithm is compatible with transformer-based architectures, as transformers’ gradient information can be used to generate UPGD perturbations. However, UCLC (zheng2022clp) for ABI is incompatible with the attention mechanism. The UCLC is designed for feature channels of convolutional layers, which do not exist in the attention mechanism. Therefore, we design a tailored novel pruning method against backdoor neurons in the attention-based models. Specifically, we consider an attention head as a “neuron”, the new pruning method then calculates the singular value decomposition of each attention head as an index of relevance to the backdoor effect. The process can also be considered as solving for the approximation of the Lipschitz Constant for each attention head according to (zheng2022clp). Then, the sensitive heads are updated using the following equation:
(4) ${\bm{\theta}}_{l}^{(k)}=\left\{\begin{aligned} &{\bm{\theta}}_{l}^{(k)}\times 0.5,&\sigma({\bm{\theta}}_{l}^{(k)})>\text{mean}(\sigma({\bm{\theta}}_{l}))+u\times\text{std}(\sigma({\bm{\theta}}_{l}))\\ &{\bm{\theta}}_{l}^{(k)},&\text{otherwise,}\end{aligned}\right$.
@@
Dataset Victim Surrogate Parameters (M) No Defense vanilla FT FT-SAM (Zhu2023ftsam) FST (min2023fst)
@@ C.4. Backdoor analysis by feature space inversion
We also provide a feature space analysis for different attacks by using a feature mask to decouple the benign and backdoor features following BTI-DBF (xu2024btidbf) and BAN (xu2024ban). The decoupling assumes that benign features related to the correct prediction introduce a lower loss, and the backdoor features related to backdoor prediction introduce a higher loss. In particular, the benign and inversed backdoor features are decoupled as follows:
@@
where $g=f_{L-1}\circ\cdots f_{1}$ is the model without the classification head. $\bm{m}$} is the learned feature mask, and $\mathcal{D}_{l}$ is a small set of benign samples with correct labels. As validated with BTI-DBF (xu2024btidbf) and BAN (xu2024ban), benign and backdoor features can be decoupled by the mask $\bm{m}$}, after which backdoor features will introduce a substantially higher loss with respect to the ground truth label.
@@
Hyperparameters for the Inversed Backdoor Feature Loss Following the settings in BTI-DBF (xu2024btidbf) and BAN (xu2024ban), we use Adam and the learning rate of 0.01 to search for 20 epochs for the feature mask in Equation 5. The optimization of the mask uses 1% of the training data. $\lambda$ is 0.72. The elements in the mask are limited to continuous values between 0 and 1.
@@
Dirty-Label
Grond
PR No Defense CLP (zheng2022clp) FT-SAM (Zhu2023ftsam)
@@
C.5. Dirty-label Grond
@@
C.6. Adversarial Backdoor Injection Does Not Impact Backdoor Effectiveness in Case of No Defense
@@ C.7. Further TAC analysis
Table 18. Evaluation of Grond against adversarial training (AT) on CIFAR10. RA refers to accuracy under PGD (madry2018pgd) adversarial attack.
Arch
No Defense AT
@@
Figure 12. The weight changes after FT-SAM (Zhu2023ftsam) for baseline attacks.
@@ C.9. Weight Changes After FT-SAM with 10 Baseline Attacks
Figure 12 provides the results of weight changes after applying FT-SAM (Zhu2023ftsam) to ten baseline attacks. The weight changes in baseline attacks show a significant difference in a few prominent neurons.
@@ C.10. Attempts to Craft Inputs that Minimize Changes in the Parameter Space
In the previous design of Grond, the backdoor is injected by the UPGD trigger. Then, Grond, through ABI, limits the changes of backdoor neurons by reducing their weights. We consider replacing the ABI component with an operation only in the training data, which does not require control over training. Our goal is to partially break the effectiveness of injecting the backdoor via the poisoned data. Specifically, we add poisoned data into the target class with a poisoning rate of $PR=5\%$ under the clean-label setting, following the standard setting in previous work (wang2023unicorn; xu2024ban). Then, we add poisoned data into every class with a poisoning rate of $PR\times\lambda$. We aim to use the poisoned data in non-targeted classes to partially break the effectiveness of poisoned data in the targeted class. Table 19 provides the experimental results with different $\lambda$ values. With the decrease in $\lambda$, both ASR (no defense) and ASR (after applying CLP) increase. This means that our attempt to minimize changes in the parameter space for stealthiness works. However, ASR after applying CLP is still lower than 10%, which means the attack is not robust against the backdoor defense.